# Multi-Atlas Brain Network Classification through Consistency Distillation and Complementary Information Fusion

### v1

@@ 5.5 Ablation Study
5.6 Time Efficiency
We conducted an experiment to compare the total runtime cost of AIDFusion with other multi-atlas baselines. The results, reported in Table 4, demonstrate that AIDFusion requires dramatically fewer epochs to converge, resulting in significantly less time spent on ABIDE, ADNI, and PPMI datasets. For the Mātai dataset, AIDFusion’s time cost is still comparable with the other baselines. Besides, since AIDFusion does not contain any repeat layers as other baselines do, it has fewer parameters and thus results in higher efficiency. This showcases the efficiency of the proposed AIDFusion.
Table 4: Time efficiency analysis. Total time (h) was recorded with a single run (including training, validation, and test) with 10-fold CV.
@@ 5.6 Time Efficiency
6 Conclusion
In this paper, we presented the Atlas-Integrated Distillation and Fusion network (AIDFusion), a novel approach to multi-atlas brain network classification. The disentangle Transformer mechanism, combined with inter-atlas message-passing and consistency constraints, effectively integrates complementary information across different atlases and ensures cross-atlas consistency at both the subject and population levels. Our extensive experiments on four fMRI datasets demonstrate that AIDFusion outperforms state-of-the-art methods in terms of classification accuracy and efficiency. Moreover, the patterns identified by AIDFusion align well with existing domain knowledge, showcasing the model’s potential for providing interpretable insights into neurological disorders. Currently, our discussion of multi-atlas brain networks is restricted to 3 atlases, i.e., AAL, Schaefer, and HO. It is also worth extending the exploration to more atlases to find out which one benefits our model the most. Besides, experiments for atlases with various numbers of ROIs (Appendix F) indicate that AIDFusion performs better when the two atlases have a similar number of ROIs. How to utilize atlases with different resolutions remains a problem. Exploring datasets with more neuroimaging modalities and more atlases with different resolutions will provide insights into how atlas resolution affects the performance of multi-atlas methods. We plan to advance multi-scale brain network analysis by creating public benchmarks and leveraging inter-scale information to capture the brain’s hierarchical organization more effectively.
Appendix A Difference between Multi-atlas and Multi-template Methods
In brain MRI analysis, the concepts of template and atlas can sometimes be confused. The atlas in our paper refers to a detailed map of brain structures [47], often derived from anatomical, functional, and histological data, and includes labels for different brain regions based on specific criteria. A template, on the other hand, is a standard reference image that serves as a common coordinate system for comparing different brain images [46]. It is usually created by averaging brain images from a group of subjects, providing a standardized space for registering or aligning individual brain images. This allows for comparison and combination of data across different subjects or groups. Figures 3 and 4 illustrate the difference between atlas and template. Although some existing works [37, 38, 30, 31] are named multi-atlas, they are more akin to multi-template methods. Instead of registering brain images to different spaces in multi-template methods, the multi-atlas methods discussed in our paper segment brain images in the common space to define ROIs differently.
Figure 3: AAL116 and Schaefer100 atlases. Each atlas is based on a different parcellation hypothesis.
Figure 4: MNI template T1-w image.
Appendix B Notation Table
Notation-wise, we use calligraphic letters to denote sets (e.g., $\mathcal{X}$), bold capital letters to denote matrices (e.g., $\boldsymbol{X}$), and strings with bold lowercase letters to represent vectors (e.g., $\boldsymbol{x}$). Subscripts and superscripts are used to distinguish between different variables or parameters, and lowercase letters denote scalars. We use $\boldsymbol{S}[i,:]$ and $\boldsymbol{S}[:,j]$ to denote the $i$-th row and $j$-th column of a matrix $\boldsymbol{S}$, respectively. Table 5 summarizes the notations used throughout the paper.
Table 5: Notation table
Notation Description
$\mathcal{D}$ Input dataset
$\boldsymbol{X}_{a},\boldsymbol{X}_{b}$ A brain network of atlases $a$ and $b$
$y_{X}$ Label of $\boldsymbol{X}_{a}$ and $\boldsymbol{X}_{b}$
$n_{a},n_{b}$ Number of nodes in atlases $a$ and $b$
$\boldsymbol{H}_{ID}$ Identity-encoded feature matrix
$\boldsymbol{m}$ Brain network representation
$d$ Dimensionality of node representations
$\boldsymbol{W}_{ID},\boldsymbol{W}_{INC},\boldsymbol{W}_{GC},\boldsymbol{W}_{Q},\boldsymbol{W}_{K},\boldsymbol{W}_{V}$ Parameter matrices
$\boldsymbol{H}$ Node representations after inter-atlas message-passing
$\boldsymbol{A}^{ab}$ The cross-atlas adjacency matrix for atlases $a$ and $b$
$\boldsymbol{H}^{ab}$ The combined node representations for atlases $a$ and $b$
$i,j$ Index for matrix dimensions
$bz$ Batch size
Appendix C An example of the Adjacency Matrix for Inter-Atlas Message-Passing
As shown in Figure 5, we generated an adjacency matrix for Schaefer100 and AAL116 by setting $k=5$. The connections between spatial neighborhoods across atlases are constructed. Note that node $v$ in atlas 1 is one of the nearest $k$ neighbors of node $u$ in atlas 2 does not mean $u$ is one of the nearest $k$ neighbors of $v$, thus the adjacency matrix is asymmetry.
Figure 5: The adjacency matrix for inter-atlas message-passing.
Appendix D Detailed Dataset Description
ABIDE The ABIDE initiative supports the research on ASD by aggregating functional brain imaging data from laboratories worldwide. ASD is characterized by stereotyped behaviors, including irritability, hyperactivity, depression, and anxiety. Subjects in the dataset are classified into two groups: TC and individuals diagnosed with ASD.
ADNI The ADNI raw images used in this paper were obtained from the ADNI database (adni.loni.usc.edu). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging (MRI), PET, other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment (MCI) and early AD. For up-to-date information, see www.adni-info.org. We include subjects from 6 different stages of AD, from cognitive normal (CN), significant memory concern (SMC), mild cognitive impairment (MCI), early MCI (EMCI), late MCI (LMCI) to AD.
PPMI The PPMI is a comprehensive study aiming to identify biological markers associated with Parkinson’s risk, onset, and progression. PPMI comprises multimodal and multi-site MRI images. The dataset consists of subjects from 4 distinct classes: normal control (NC), scans without evidence of dopaminergic deficit (SWEDD), prodromal, and PD.
Mātai Mātai is a longitudinal single site, single scanner study designed for detecting subtle changes in the brain due to a season of playing contact sports. This dataset consists of the brain networks preprocessed from the data collected from Gisborne-Tairāwhiti area, New Zealand, with 35 contact sport players imaged at pre-season (N=35) and post-season (N=25) with subtle brain changes confirmed using diffusion imaging study due to playing contact sports.
Appendix E Implementation Details
The settings of our experiments mainly follow those in [16]. We split each dataset into 8:1:1 for training, validation and test, respectively. We evaluate each model with the same random seed under 10-fold cross-validation and report the average accuracy. The whole network is trained in an end-to-end manner using the Adam optimizer [21]. We use the early stopping criterion, i.e., we halve the learning rate when there is no further improvement on the validation loss during 25 epochs and stop the training once the learning rate is smaller than the minimum rate we set. All the codes were implemented using PyTorch [40] and Deep Graph Library [50] packages. All experiments were conducted on a Linux server with an AMD Ryzen Threadripper PRO 5995WX 64-Cores and an NVIDIA GeForce RTX 4090. The version for the software we used in AIDFusion is listed in Table 6.
Table 6: The software dependency of AIDFusion.
Dependency Version
Python 3.10.13
cudatoolkit 12.2
pytorch 2.2.1+cu121
DGL 2.1.0+cu121
scikit-learn 1.4.1.post1
numpy 1.26.4
matplotlib 3.8.3
nilearn 0.10.4
All the baselines used in this paper are implemented by ourselves. In AIDFusion, we adopt a mean pooling layer as the readout function and a two-layer MLP with ReLU as the prediction head. The number of incompatible nodes $r$ is set to 4. The number of clusters $n^{\prime}$ for the cluster assignment matrix $\boldsymbol{S}$ is set to half of the average number of nodes for all atlases, e.g., for Schaefer100 and AAL116, $n^{\prime}=(100+116)/2/2=54$. The temperature hyper-parameter $\tau$ in Eq. (10) is set to 0.75. The hidden dimensions of all layers are set to 100. The batch size of each dataset is set to 10% of the subject number in the dataset. We tuned the other hyperparameters on validation set including $k$ in $k$NN for the inter-atlas message-passing, trade-off hyperparameters $\lambda_{1},\lambda_{2},\lambda_{4},\lambda_{4}$ for loss function Eq. (12), the initial learning rate $init\_lr$, and the minimum learning rate $min\_lr$ used for early stop. To be specific, we search $k$ from {3, 5, 10}, $\lambda_{1}$ from {1e-2, 1e-1, 1e0, 1e1, 1e2}, $\lambda_{2}$ from {1e-2, 1e-1, 1e0, 1e1, 1e2}, $\lambda_{3}$ from {1e-6, 1e-5, 1e-4, 1e-3, 1e-2}, $\lambda_{4}$ from {1e-1, 1e0, 1e1}, $inir\_lr$ from {5e-5, 8e-5, 1e-4, 2e-4, 7e-4} and $inir\_lr$ from {6e-5, 1e-5, 5e-6, 1e-6}. The optimized hyperparameters for AIDFusion is reported in Table 7.
Table 7: The optimized hyperparameters for AIDFusion.
ABIDE ADNI PPMI Mātai
$k$ 10 5 5 5
$\lambda_{1}$ 1e1 1e1 1e1 1e-2
$\lambda_{2}$ 1e-1 1e1 1e1 1e0
$\lambda_{3}$ 1e-2 1e-5 1e-4 1e-3
$\lambda_{4}$ 1e0 1e0 1e0 1e0
$init\_lr$ 1e-4 8e-5 1e-4 7e-4
$min\_lr$ 6e-5 6e-5 6e-5 6e-5
Appendix F Results of More Evaluation Metrics
Table 8: Results of more evaluation metrics on ABIDE dataset. The best result is highlighted in bold.
Precision Recall micro-F1 ROC-AUC
MGRL 59.90 ± 6.42 60.03 ± 6.31 59.71 ± 5.10 61.39 ± 5.08
MGT 60.33 ± 4.78 69.29 ± 6.74 64.21 ± 3.62 63.60 ± 3.80
METAFormer 59.33 ± 4.05 61.91 ± 7.79 60.20 ± 4.26 61.31 ± 3.94
LeeNet 64.30 ± 5.24 43.04 ± 6.09 51.23 ± 4.74 60.44 ± 3.06
AIDFusion (ours) 62.25 ± 3.00 74.80 ± 4.38 67.90 ± 3.12 66.73 ± 3.25
In addition to accuracy, we also report other evaluation metrics, including precision, recall, micro-F1, and ROC-AUC, for all the multi-atlas deep models on the ABIDE dataset. As displayed in Table 8, AIDFusion performs the best across all these metrics except for precision. We observe that compared to other baselines, our AIDFusion can significantly improve recall without compromising precision. Moreover, in medical diagnostics, it is crucial to ensure that all individuals with a certain condition are correctly identified, even if it leads to some false positives. Missing a true positive (failing to diagnose a disease) can have severe consequences, while false positives can be further examined or retested. Therefore, models with higher recall rates, such as our AIDFusion, are more suitable for real-life medical auxiliary diagnosis.
Table 9: Results of more atlases on ADNI dataset. The best results for each atlas setting are highlighted in bold.
Atlas model acc ± std
Schaefer100 AAL116 HO48
MGRL 62.74 ± 3.55
MGT 63.99 ± 4.34
METAFormer 66.52 ± 2.63
LeeNet 64.63 ± 1.34
AIDFusion 67.57 ± 2.04
MGRL 64.48 ± 1.68
MGT 60.48 ± 1.91
METAFormer 64.48 ± 1.68
LeeNet 64.48 ± 1.58
AIDFusion 65.99 ± 2.62
MGRL 57.92 ± 2.82
MGT 60.64 ± 3.52
METAFormer 65.23 ± 3.25
LeeNet 63.42 ± 1.82
AIDFusion 65.91 ± 1.80
MGRL 56.63 ± 4.66
MGT 62.89 ± 1.28
METAFormer 66.33 ± 2.80
LeeNet 64.40 ± 1.71
AIDFusion 66.59 ± 1.77
Appendix G Results with More Atlases
To further evaluate the effectiveness of AIDFusion with other atlases, we conducted experiments using an additional atlas, HO48 [36], on the ADNI dataset. The results, presented in Table 9 indicate that the proposed AIDFusion achieves the best performance across all four atlas settings. Notably, increasing the number of atlases does not necessarily enhance model performance. In some cases, using all three atlases yields lower accuracy compared to the combination of Schaefer100 and AAL116. Additionally, the choice of atlas combination is crucial for multi-atlas methods. In dual-atlas experiments, combining two atlases with a similar number of ROIs (Schaefer100 and AAL116) can mutually enhance and significantly improve model performance.
To further explore how the resolution of atlases will influence the performance of our model, we conduct experiments for atlases with various numbers of ROIs. The Schaefer atlas allows adjusting the resolution of ROIs (e.g., from 100 to 1000). We selected Schaefer100 for detailed study because a previous study [54] found that using 100 ROIs with the Schaefer atlas usually performs better than using more ROIs. To verify this conclusion in multi-atlas brain network classification, we conducted experiments using AAL116 combined with Schaefer200, Schaefer500, and Schaefer1000. Results showed that AAL116 combined with Schaefer100 achieves the best results. Additionally, we provided experimental results using the HO48 atlas in Appendix F. The dual-atlas experiments demonstrate that combining two atlases with a similar number of ROIs (AAL116 and Schaefer100) can mutually enhance and significantly improve model performance. It is also interesting for us to explore using both atlases of around 200 nodes (or around 1000 nodes). We will leave such exploration about multi-scale brain networks in the future.
Table 10: Results of more atlases with different resolutions on ABIDE dataset. The best results for each atlas setting are highlighted in bold.
Atlas 1 Atlas 2 acc ± std
AAL116 Schaefer100 66.35 ± 3.26
AAL116 Schaefer200 65.03 ± 5.10
AAL116 Schaefer500 64.72 ± 4.80
AAL116 Schaefer1000 63.15 ± 2.80
Figure 6: Visualization for attention maps on ABIDE. VIS = visual network; SMN = somatomotor network; DAN = dorsal attention network; VAN = ventral attention network; LN = limbic network; FPCN = frontoparietal control network; DMN = default mode network.
Appendix H Brain Network Attention Map on ABIDE
In the ASD analysis using the ABIDE dataset (Figure 6), AIDFusion identifies common connections within the lingual gyrus of the VIS network in both Schaefer100 and AAL116. This aligns with existing ASD studies which suggest a greater reliance on visual perceptual processing and more effortful top-down control during semantic processing in ASD [45]. Besides, for Schaefer, AIDFusion emphasizes ROIs in the DAN, particularly the connection between the right posterior and VIS, consistent with Koshino et al. [23] who found lower activation in ASD subjects in the inferior left prefrontal area (verbal processing and working memory) and the right posterior temporal area (theory of mind processing). In the AAL analysis, connections in the DMN and the frontoparietal control network (FPCN) are highlighted, supporting the understanding that (1) dysfunctions in DMN nodes and their interactions contribute to difficulties of ASD in attending to socially relevant stimuli [39], and (2) the ASD group shows reduced lateral frontal activity and diminished hippocampal connectivity, especially between the hippocampus and FPCN regions [8]. These findings clarify why the features identified by AIDFusion are distinctive for ASD biomarker.
Appendix I Ablation Study for Incompatible Nodes
In this section, we further explore the function of incompatible nodes by visualizing the attention map of ADFusion w/o incompatible nodes. The attention maps are shown in Figure 7. We can observe that, when not using incompatible nodes, the attentions of two atlases (in the right column) are remarkably imbalanced. Attentions on Schaefer are much higher than those on AAL. Besides, the attention map of AAL exhibits over-smoothing and no highlighted network is found, which indicates the model is not able to extract the distinguishable connections. This case study demonstrates that the incompatible nodes enable the model to filter out the inconsistent atlas-specific information.
Appendix J Hyperparameter Analysis
In this section, we study the sensitivity of four trade-off hyperparameters in Eq. (12). All experiments are conducted on the ADNI dataset. We tune the value of $\lambda_{1}$ from 1e0 to 1e2, $\lambda_{2}$ from 1e0 to 1e2, $\lambda_{3}$ from 1e-6 to 1e-4 and $\lambda_{4}$ from 1e-1 to 1e1. The results presented in Table 11 show that our model performs the best when $\lambda_{1}=$ 1e1, $\lambda_{2}=$ 1e1, $\lambda_{3}=$ 1e-5 and $\lambda_{4}=$ 1e0. We can exhibit that these trade-off hyperparameters in the loss function will marginally affect the model performance on ADNI (less than 1%), which demonstrates the stability of AIDFusion.
Table 11: The hyperparameter sensitivity analysis for AIDFusion on ADNI dataset.

### v2

@@ V-D Ablation Study
V-E Model Interpretation
Biomarker Detection. In neurodegenerative disorder diagnosis, identifying salient ROIs/connections associated with predictions as potential biomarkers is crucial. In this study, we utilize attention scores from the Transformer layer to generate heat maps for brain networks to interpret our model. We visualize these attention maps using the Nilearn toolbox [63]. Figure 2 presents attention maps for two atlases, where higher attention values mean better classification potential for AD (from the ADNI dataset). We utilized 7 networks [64] to assess the connections between our highlighted ROIs and major networks potentially involved with disorders. ROIs from the AAL that do not overlap with these seven networks are excluded from the heat maps. The top 10 ROIs with the highest attention values are displayed in the brain view. As depicted in the attention maps, attention maps of both Schaefer and AAL atlases identify common connections between the visual network (VIS) and the dorsal attention network (DAN), recognized as key connectivities in AD research [65, 66]. Additionally, atlas-specific connections are highlighted. For example, the attention map of Schaefer atlas emphasizes connections within the default mode network (DMN) corresponding with the observations of Damoiseaux et al. [67]. Findings on the attention map on AAL are consistent with Agosta et al. [66], showing that AD is associated with connectivities in VIS, especially in frontal networks. These findings suggest that AIDFusion effectively captures complementary information from different atlases. Besides, we also find some of our findings diverge from conventional neuroscientific understanding. For example, connections between VIS and somatomotor network (SMN) have a high attention weight in AIDFusion on Schaefer atlas, which may imply AD is related to the function of defining the targets of actions and providing feedback for visual activation. This insight has not been identified by existing literature in the domain of neuroscience. We also note that the attention map over AAL116 reveals that the cerebellar regions (the lower portion of the VIS network) receive noticeably lower attention weights compared to cortical ROIs. This observation suggests that AIDFusion is capable of downweighting atlas-specific regions that lack consistent correspondence across atlases (e.g., cerebellum absent in Schaefer100), thereby preventing spurious feature alignment. It highlights the model’s ability to selectively focus on semantically meaningful and cross-atlas-consistent regions during representation learning.
Case Study for Incompatible Nodes. To further support the function of incompatible nodes, we conducted a case study comparing attention maps of AIDFusion with and without the incompatible nodes. The attention maps are shown in Figure 3. We can observe that, when not using incompatible nodes, the attentions of two atlases (in the right column) are remarkably imbalanced. Attention on Schaefer is much higher than that on AAL. Besides, the attention map of AAL exhibits over-smoothing and no highlighted network is found, which indicates the model is not able to extract the distinguishable connections. This case study demonstrates that the incompatible nodes enable the model to filter out the inconsistent atlas-specific information.
TABLE VII: Brain Network Classification Results (Average Accuracy ± Standard Deviation) over 10-fold-CV under out-of-distribution setting. The best results are highlighted in bold.
model Accuracy Precision Recall micro-F1 ROC-AUC
MGRL 62.53 ± 4.20 60.37 ± 4.52 62.27 ± 7.60 61.13 ± 5.25 62.49 ± 4.25
MGT 56.78 ± 3.99 55.69 ± 5.08 49.40 ± 8.63 51.73 ± 6.17 56.43 ± 3.93
METAFormer 61.17 ± 2.92 58.71 ± 3.21 62.70 ± 8.13 60.37 ± 4.48 61.23 ± 3.00
LeeNet 59.51 ± 3.61 62.06 ± 5.20 39.36 ± 15.32 46.21 ± 12.83 58.58 ± 4.05
AIDFusion (ours) 63.50 ± 3.72 59.67 ± 3.22 72.55 ± 6.13 65.37 ± 3.69 63.91 ± 3.74
TABLE VIII: Time efficiency analysis. Total time (h) was recorded with a single run (including training, validation, and test) with 10-fold-CV.
@@ V-E Model Interpretation
V-F Generalization Ability
While task-specific biomarkers are valuable for identifying disease-relevant features, it is crucial to determine whether these biomarkers are invariant over the entire population, i.e., whether they generalize well across sub-populations and other diverse populations. To assess the generalization ability of AIDFusion, we conduct evaluations on subjects from previously unseen sites. Specifically, we have conducted additional experiments under an out-of-domain (OOD) setting following the protocol proposed in [68], where one site is held out as the OOD test set for each fold. This simulates the realistic scenario of deploying the model to an unseen clinical center.
We report the performance of AIDFusion and all multi-atlas baselines on ABIDE dataset, in which the test set includes both in-domain (ID) and OOD subjects. As shown in Table VII, AIDFusion consistently achieves superior performance compared to other multi-atlas methods. This demonstrates the robustness and transferability of our model across sites with heterogeneous data distributions.
V-G Time Efficiency
We conducted an experiment to compare the total runtime cost of AIDFusion with other multi-atlas baselines. Note that our method requires computing population-level consistency, which scales with the number of subjects rather than being a linear per-subject cost. The results, reported in Table VIII, demonstrate that AIDFusion requires dramatically fewer epochs to converge, resulting in significantly less time spent on ABIDE, ADNI, and PPMI datasets. For the Mātai dataset, AIDFusion’s time cost is still comparable with the other baselines. Besides, since AIDFusion does not contain any repeat layers as other baselines do, it has fewer parameters and thus results in higher efficiency. This showcases the efficiency of the proposed AIDFusion.
V-H Hyperparameter Analysis
We study the sensitivity of four trade-off hyperparameters in Eq. (13). All experiments are conducted on the ADNI dataset. We tune the value of $\lambda_{1}$ from 1e0 to 1e2, $\lambda_{2}$ from 1e0 to 1e2, $\lambda_{3}$ from 1e-6 to 1e-4 and $\lambda_{4}$ from 1e-1 to 1e1. The results presented in Table IX show that our model performs the best when $\lambda_{1}=$ 1e1, $\lambda_{2}=$ 1e1, $\lambda_{3}=$ 1e-5 and $\lambda_{4}=$ 1e0. We can exhibit that these trade-off hyperparameters in the loss function will marginally affect the model performance on ADNI (less than 1%), which demonstrates the stability of AIDFusion.
TABLE IX: The hyperparameter sensitivity analysis for AIDFusion on ADNI dataset over 10-fold-CV of a single run.
@@ V-H Hyperparameter Analysis
VI Conclusion
In this paper, we presented the Atlas-Integrated Distillation and Fusion network (AIDFusion), a novel approach to multi-atlas brain network classification. The disentangle Transformer mechanism, combined with inter-atlas message-passing and consistency constraints, effectively integrates complementary information across different atlases and ensures cross-atlas consistency at both the subject and population levels. Our extensive experiments on four fMRI datasets demonstrate that AIDFusion outperforms state-of-the-art methods in terms of classification accuracy and efficiency. Moreover, the patterns identified by AIDFusion align well with existing domain knowledge, showcasing the model’s potential for providing interpretable insights into neurological disorders. We hope our work inspires further research in multi-atlas brain network analysis and demonstrates its significance in real-world applications, such as early diagnosis and personalized treatments for neurodegenerative diseases.
Acknowledgments
This research/project is supported by the Ministry of Education, Singapore under its Academic Research Fund Tier 1 (RG16/24). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.